# Mechanistic Indicators of Understanding in Large Language Models

### v4

@@ 3.2 Internal Models of External States: The Othello-GPT Findings
Nanda et al. (2023b) trained 64 distinct linear probes, one for each square of the Othello board. Each probe was designed to predict the state of its assigned square by examining the model’s activations in the last token’s residual stream, at the final step before the next legal move prediction (see Fig. 8b). Since each square has three possible states—it can be occupied by a black disc, a white disc, or remain empty—they trained each probe to learn three distinct weight vectors within the activation space, one for each of the possible square states (in this sense, each probe can be viewed as a combination of three distinct probes, as a single probe usually targets just one feature). In contrast to earlier studies that sought and failed to find linear encodings of “Black,” “White,” and “Empty” features, however, the key idea of Nanda et al. was to look for features that are relative to the perspective of the current turn’s player: “Mine,” “Yours,” and “Empty.” These features correspond to three directions in the residual stream’s latent space—the “Mine,” “Yours,” and “Empty” directions. To predict a square’s status, the probe projected the activation vector onto these three directions and identified the one yielding the highest score, thereby determining the most likely state. [^23] A softmax function ensures the three projections sum to 1, converting them into probabilities.
@@
Yet when Li et al. applied principal component analysis (PCA) to the directions identified by their probes—projecting them into a 3D space to reveal the three dimensions capturing the most variance—they found a spatial arrangement resembling the actual Othello board layout (2023, Appendix). We replicated this analysis using Nanda et al.’s linear probes and two dimensions, and found an even more pronounced correspondence: the directions picked out by the probes to represent spatially adjacent squares are themselves adjacent in the projection, with rows forming roughly linear arrangements and columns doing likewise, reproducing the grid structure of the board itself (see Fig. 9a). [^25] Some caveats apply. First, we do not carry out the last step of the typical MI pipeline: ensuring the identified signal is causally relied on by the model. Doing so would address a methodological concern raised by Li et al. (2023), who observed some board structure—though considerably weaker and more jumbled—even when applying PCA to probes trained on random linear projections of Othello moves, suggesting spatial signal might partly arise from the probing method itself. Second, the model might encode and exploit spatial structure in many different ways internally: Yuan and Søgaard (2025), for example, found that token embeddings for Othello moves (e.g., “D4”) also cluster by spatial proximity. What we can conclude with reasonable certainty is that structural isomorphism is present in some form within the model’s representations; demonstrating precisely how it is causally exploited awaits further investigation.
@@
Third, there is the argument from causal efficacy: the board representation clearly plays a causal role in the move predictions. In line with philosophical accounts of representation in LLMs (Yetman, 2025; Williams, 2026), both teams of researchers went beyond observing correlations and demonstrated causal role via interventions. By manually “flipping” the representation of a disc from “Black” to “White” (Li et al., 2023) or erasing a previously played disc (Nanda et al., 2023b), they were able to get the model to output moves that were legal only in the counterfactual board states and not in the original board states. This satisfies a robust definition of causal role: changing the internal representation systematically and predictably alters the model’s behavior. This shows that the model is not just predicting the next token based on previous tokens, but relying on its internal representation of the board.
@@
And crucially, Othello-GPT’s state-of-the-world understanding is relational and integrated. It does not treat each square state as an isolated fact, but constructs a representation of the board that enables it to attend to these facts together, as a system of facts. Evidence of this emerged when Li et al. (2023) used “latent saliency maps” to visualize which squares played the largest role in informing the model’s top next-move prediction. Each square’s contribution was measured by how much changing the internal representation of that square state lowered the prediction’s likelihood. They did this with two versions of Othello-GPT: one trained only on computer-generated sequences of random moves (synthetic data), the other trained on expert human games (championship data). While the version trained on synthetic data showed high saliency only for the squares that were strictly necessary to rendering a move legal, the version trained on championship data exhibited a much broader saliency pattern, mirroring how the moves of expert players depend on global features of the board (see Fig. 9b). Such a grasp of how elements of a system depend on each other is considered a hallmark of understanding (Riggs, 2003; Grimm, 2011, 2025).
@@
Moreover, the fidelity of this emergent state-of-the-world understanding is not uniform. Nanda et al. observed “end-game degradation,” where the model’s reliance on its internal representation of the board state seemed to falter in the final moves of a game. They hypothesized that as the board fills up and options become fewer, it becomes computationally cheaper for the model to switch from a complex world-modeling circuit to simpler heuristics. This suggests that LLMs may not be consistent state-of-the-world trackers. They may pragmatically toggle between modeling the state of the world and relying on shallow heuristics, depending on which circuit minimizes loss most efficiently. Thus, while the Othello-GPT findings suggest that LLMs can maintain an internal representation of the state of the external world, they also warn us that they will not necessarily do so when a shortcut is available.
@@ 4.1 A Simple Circuit: The Induction Head
A simple example of a circuit is the induction head, first identified in a two-layer toy model by Elhage et al. (2021). This circuit provides a mechanism for a basic but crucial form of in-context learning: completing repeating patterns. Specifically, it enables a model to solve tasks of the form “[A] [B] … [A] → [B].” This is an extremely useful pattern for natural language, as it allows the model to correctly complete a name like “Michael Jordan” by recognizing that this instance of “Michael” is the one that was previously followed by “Jordan” in the same text, rather than “Michael Jackson” or “Michael Faraday.”
@@
  • Step 2: Retrieving the Completion. Later in the model, a second attention head—the induction head proper—processes the second instance of token [A]. Its query essentially asks: “Is there a token earlier in the text that was marked as ‘coming after a token like me’?” It scans the keys of all previous tokens and finds a strong match with the first token [B], which was marked by the prefix head. The induction head then copies the relevant information from [B]’s residual stream into the current stream, strongly upweighting the logit for [B] as the next token. [^26] Remember from section 2 that the information gating is governed by key-query matching. The first attention head at layer one does this matching using positional encodings: it matches a query “I’m looking for position x-1” with the key “I’m at position x-1.” The second attention head at layer two matches the query “I’m looking for a token after an [A]” with the key “I’m after an [A].” This allows information to flow in the desired way. For a great visual explanation, we recommend (McDougall, 2023).
While this learned circuit still falls short of embodying a principle connecting facts, it constitutes an important first step towards it, because it implements a content-agnostic algorithm. [^27] On the importance of non-content-specific representations in neural networks, see Shea (2021). The mechanism works just as well for completing “Harry Potter” as it does for “Marie Curie” or for repeating arbitrary sequences of random tokens. It is not learning a content-specific statistical association between “Michael” and “Jordan”; rather, it is implementing a general procedure for sequence completion by looking back in a targeted way.
The initial evidence for induction heads came from toy models. But subsequent behavioral analyses suggest that they are also a crucial component of in-context learning in large-scale models “in the wild” (Olsson et al., 2022). [^28] They also study induction heads responsible for “fuzzy” copying tasks of the form [A*] [B*] … [A] → [B], such as in translation where [A*] and [B*] correspond to [A] and [B] in another language. Along those lines, Feucht et al. (2025) recently proposed and empirically investigated a distinction between token induction heads and concept induction heads. While their evidence is primarily behavioral (see Variengien, 2023), it clearly emerges that models are able to look back in the text in targeted ways to recall relevant token sequences. The induction head thus serves as a compelling first piece of evidence that LLMs can unify many specific instances of pattern completion under a single, learned circuit.
@@ 4.2 A Deeper Principle: The Fourier Algorithm for Modular Addition
While the induction head circuit does not connect facts or embody principled understanding, MI research has also found a striking example of a circuit that does seem to embody principled understanding. Nanda et al. (2023a) trained a simple one-layer transformer exclusively on examples of modular addition. Modular addition is addition on a circle or loop instead of a straight number line. This ensures that the numbers are always kept within a specific range, as they are with the hours on a clock or the days of the week. To perform modular addition, one adds the numbers normally before applying the modulus (the size of the loop): one calculates how many times the modulus fits into the sum of the numbers and takes the remainder. This remainder is the solution. When adding 9 + 5 with a modulus of 12, for example, one first calculates 9 + 5 = 14 before seeing how many times 12 goes into 14. Since 12 goes into 14 only once and leaves a remainder of 2, the solution to 9 + 5 (modulo 12) is 2. The model was trained on examples of modular addition using a modulus of 113 (e.g. 55 + 71 = 13), but was never given the rules of addition or the definition of the modulo operator. If LLMs were simply n-gram models, one would expect the model to approach this task by simply memorizing the entire addition table given in the training data.
@@
This simple circuit functionally satisfies the main conditions for explanatory understanding discussed in contemporary epistemology (see Baumberger et al., 2017). It embodies a network of explanatory connections (Khalifa, 2017; Strevens, 2013), which satisfy the condition of factivity (Kvanvig, 2009; Kelp, 2015), since they rely on geometric facts (the trigonometric identities) that actually obtain. These connections enable the unification of many mathematical facts under a single principle (Friedman, 1974; Kitcher, 1981), which in turn allows the model to compress its repository of memorized cases down to a single circuit implementing this principle (Wilkenfeld, 2018). This also bestows upon the model the inferential capacities associated with grasping (Hills, 2015; Grimm, 2011)—manifested here in the model’s ability to generalize correctly to all inputs, not just those in the training set. Finally, the circuit satisfies multiple criteria linked to the condition of justification: it has a wide scope, is used consistently and reflects a pursuit of simplicity via weight decay (Baumberger et al., 2017; Beckmann, 2025).
@@ 4.3 Circuits in the Wild
While the modular addition circuit was found in a toy model, a growing body of research is uncovering similarly complex circuits in state-of-the-art LLMs. Research on regular (non-modular) addition “in the wild” suggests that real LLMs also rely on operations across multiple frequencies—or, equivalently, across multiple moduli (Zhou et al., 2024; Ameisen et al., 2025; Kantamneni and Tegmark, 2025; Nikankin et al., 2025). Other examples of circuits in the wild include an indirect object identification circuit (Wang et al., 2023), a greater-than circuit for numerical comparison (Hanna et al., 2023), and a multiple-choice question circuit (Lieberum et al., 2023).
Recently, researchers at Anthropic have developed techniques to automatically extract functional circuits for specific prompts from frontier models like Claude 3.5 Sonnet (Ameisen et al., 2025; Lindsey et al., 2025). They have successfully identified circuits responsible for tasks as diverse as entity recognition, planning the structure of a poem, generating medical diagnoses, and even refusing to answer unsafe queries. A particularly elegant example is a line-breaking circuit in Claude 3.5 Haiku that uses geometric transformations—rotating representations of character count relative to line length in its attention blocks—to detect when the model is approaching the end of a line (Gurnee et al., 2025). These findings confirm that the principle of computation via specialized circuits is not an artifact of toy models, but a core operating principle of modern LLMs.
@@
The primary testing ground for this capability is the Abstraction and Reasoning Corpus (ARC-AGI), which was expressly designed to test a system’s ability to discover the rules or principles underlying something at inference time, i.e. after training has been completed (Chollet, 2019; Chollet et al., 2025). To succeed on ARC, a model must be able to look at a few pairs of inputs and outputs, infer the abstract rule by which one is transformed into the other (e.g., “gravity pulls all blue pixels down until they hit a red pixel”), and apply that rule to a test input.
While humans find these tasks intuitive, LLMs typically have a harder time discovering principles in real time. [^31] A related limitation emerges from recent investigations into LLMs’ maze-solving capabilities (Palod et al., 2025; Valmeekam et al., 2025). This tendency to default to static, training-induced mechanisms has led critics to argue that LLMs lack true intelligence, reducing them to combinators of “heuristics” (Mitchell, 2025), “program templates” (Chollet, 2024), or pieces of “procedural knowledge” (Kambhampati, 2025).
@@ 5  A Motley Mix of Mechanisms
But does the existence of these capacities mean that LLMs actually understand in typical cases? Zooming out from the most compelling, fine-grained findings of mechanistic interpretability to date reveals a more complex reality. Insofar as LLMs instantiate forms of cognition, they instantiate highly strange forms of it, shaped by constraints and incentives that differ starkly from those shaping human intelligence (Fleuret, 2023; Shanahan, Murray, 2024). This divergence complicates the attribution of understanding.
@@
At inference, these mechanisms operate in parallel: LLMs typically tackle problems by deploying multiple distinct mechanisms simultaneously rather than relying on a single circuit (Lindsey et al., 2025). These mechanisms may cooperate, compete, or contribute independent pieces, with the final output emerging from their combined effect. Higher-quality circuits are sometimes reinforced, but sometimes also drowned out by lower-quality circuits.
@@
The constructive interference observed in modular addition offers a glimpse of this: the model’s final answer arises from the aggregation of five separate frequency-based calculations. Similarly, internal circuitry for standard addition exemplifies the deployment of numerous independent heuristics, such as multiple modulo mechanisms—e.g., computing a + b modulo 2 to determine if the sum is even or odd, and modulo 10 to determine its final digit (Zhou et al., 2024; Ameisen et al., 2025; Kantamneni and Tegmark, 2025; Nikankin et al., 2025). These are supplemented by parallel range-finding mechanisms of varying precision, which determine the approximate magnitude of the sum (Ameisen et al., 2025; Nikankin et al., 2025). The final answer emerges from the collective consensus of this “bag of tricks.” [^32] To compute 36 + 59, the LLM Claude 3.5 relies on the additive combination of a range-finding mechanism placing the sum at $\sim 92$; a modulo 100 mechanism according to which the sum ends with 95; and a modulo 10 mechanism according to which the sum ends with 5. Together, these mechanisms constructively interfere at 36 + 59 = 95 (Ameisen et al., 2025). Interestingly, the modulo mechanisms resemble how humans carry over digits when adding—effectively adding in steps of modulo 10, 100, 1000, and so on.
@@
Othello-GPT likewise deploys a mix of mechanisms in parallel: localized decision rules—like one that updates specific squares only when A4 is played with certain neighbors occupied; specific board pattern detectors that look for states that make moves legal; and timing-based heuristics that rely on “clock neurons” to track the game’s progress and upweight moves that are statistically more likely in the late game, such as those targeting edge squares (Lin et al., 2024).
@@
Even factual recall relies on distinct pathways operating in parallel. When prompted with “Michael Jordan plays the sport of,” Claude 3.5 uses at least two pathways. One pathway, activated by words like “plays” and “sport,” broadly boosts sports-related features, including both “basketball” and “football.” Simultaneously, a second pathway, triggered by “Michael Jordan,” strongly activates “basketball” while suppressing alternatives. The correct output—“basketball”—emerges from the confluence of these separate streams (Ameisen et al., 2025; Chughtai et al., 2024).
@@
The strangeness of LLM cognition is compounded by the drastic variation in the quality of these mechanisms. They form a “motley mix” spanning the full hierarchy of mechanisms: from pure Blockhead-style memorization of training data (Carlini et al., 2023; Aerni et al., 2024; Stoehr et al., 2024) through shortcuts that rely on spurious correlations (Geirhos et al., 2020) and shallow heuristics that rely on superficial statistics (e.g., Ma et al., 2025) to mechanisms encoding the kinds of conceptual, state-of-the-world, and principled understanding we have detailed.
@@
Beyond examining particular examples, this motley mix phenomenon can be established systematically through loss curvature analysis (Merullo et al., 2025). This approach asks a simple question: which parts of a model are such that even small weight changes cause large increases in average loss across different inputs, and which parts are such that even large changes barely affect average loss? This can be measured by examining the curvature of the loss landscape with respect to weights, averaged across many inputs. Weights implementing mechanisms recruited across many inputs reinforce the increase in loss and appear as high-curvature; weights recruited only in specific cases produce sharp increases in loss only in those cases, which wash out when averaged, registering as low-curvature or “flat.”
@@
An LLM is therefore a motley mix of mechanisms: a system where many circuits of varying quality crystallize during training and operate in parallel at inference. These mechanisms occasionally interact via activation or inhibition, but seem to lack centralized or hierarchical control. [^33] Adding a form of hierarchical control to transformer-based models is currently a hot topic in AI (Jolicoeur-Martineau, 2025; Behrouz et al., 2025). Note that the meta-cognitive features discussed in subsection 2.1 represent early signs of hierarchical processing, but function more as scattered and isolated checks. LLM outputs accordingly resemble the verdict of a gigantic committee of drastically varying expertise, working largely in isolation with only occasional collaboration.

### v5

@@ 3.2 Internal Models of External States: The Othello-GPT Findings
Nanda et al. () trained 64 distinct linear probes, one for each square of the Othello board. Each probe was designed to predict the state of its assigned square by examining the model’s activations in the last token’s residual stream, at the final step before the next legal move prediction (see Fig. 8b). Since each square has three possible states—it can be occupied by a black disc, a white disc, or remain empty—they trained each probe to learn three distinct weight vectors within the activation space, one for each of the possible square states (in this sense, each probe can be viewed as a combination of three distinct probes, as a single probe usually targets just one feature). In contrast to earlier studies that sought and failed to find linear encodings of “Black,” “White,” and “Empty” features, however, the key idea of Nanda et al. was to look for features that are relative to the perspective of the current turn’s player: “Mine,” “Yours,” and “Empty.” These features correspond to three directions in the residual stream’s latent space—the “Mine,” “Yours,” and “Empty” directions. To predict a square’s status, the probe projected the activation vector onto these three directions and identified the one yielding the highest score, thereby determining the most likely state. [^23] A softmax function ensures the three projections sum to 1, converting them into probabilities.
@@
Yet when Li et al. applied principal component analysis (PCA) to the directions identified by their probes—projecting them into a 3D space to reveal the three dimensions capturing the most variance—they found a spatial arrangement resembling the actual Othello board layout (2023, Appendix). We replicated this analysis using Nanda et al., 2023b’s linear probes and two dimensions, and found an even more pronounced correspondence: the directions picked out by the probes to represent spatially adjacent squares are themselves adjacent in the projection, with rows forming roughly linear arrangements and columns doing likewise, reproducing the grid structure of the board itself (see Fig. 9a). [^25] Some caveats apply. First, we do not carry out the last step of the typical MI pipeline: ensuring the identified signal is causally relied on by the model. Doing so would address a methodological concern raised by Li et al. (2023), who observed some board structure—though considerably weaker and more jumbled—even when applying PCA to probes trained on random linear projections of Othello moves, suggesting spatial signal might partly arise from the probing method itself. Second, the model might encode and exploit spatial structure in many different ways internally: Yuan and Søgaard (2025), for example, found that token embeddings for Othello moves (e.g., “D4”) also cluster by spatial proximity. What we can conclude with reasonable certainty is that structural isomorphism is present in some form within the model’s representations; demonstrating precisely how it is causally exploited awaits further investigation.
@@
Third, there is the argument from causal efficacy: the board representation clearly plays a causal role in the move predictions. In line with philosophical accounts of representation in LLMs (Yetman,, 2025; Williams,, 2026), both teams of researchers went beyond observing correlations and demonstrated causal role via interventions. By manually “flipping” the representation of a disc from “Black” to “White” (Li et al.,, 2023) or erasing a previously played disc (Nanda et al., 2023b, ), they were able to get the model to output moves that were legal only in the counterfactual board states and not in the original board states. This satisfies a robust definition of causal role: changing the internal representation systematically and predictably alters the model’s behavior. This shows that the model is not just predicting the next token based on previous tokens, but relying on its internal representation of the board.
@@
And crucially, Othello-GPT’s state-of-the-world understanding is relational and integrated. It does not treat each square state as an isolated fact, but constructs a representation of the board that enables it to attend to these facts together, as a system of facts. Evidence of this emerged when Li et al. (2023) used “latent saliency maps” to visualize which squares played the largest role in informing the model’s top next-move prediction. Each square’s contribution was measured by how much changing the internal representation of that square state lowered the prediction’s likelihood. They did this with two versions of Othello-GPT: one trained only on computer-generated sequences of random moves (synthetic data), the other trained on expert human games (championship data). While the version trained on synthetic data showed high saliency only for the squares that were strictly necessary to rendering a move legal, the version trained on championship data exhibited a much broader saliency pattern, mirroring how the moves of expert players depend on global features of the board (see Fig. 9b). Such a grasp of how elements of a system depend on each other is considered a hallmark of understanding (Riggs,, 2003; Grimm,, 2011, 2025).
@@
Moreover, the fidelity of this emergent state-of-the-world understanding is not uniform. Nanda et al., 2023b observed “end-game degradation,” where the model’s reliance on its internal representation of the board state seemed to falter in the final moves of a game. They hypothesized that as the board fills up and options become fewer, it becomes computationally cheaper for the model to switch from a complex world-modeling circuit to simpler heuristics. This suggests that LLMs may not be consistent state-of-the-world trackers. They may pragmatically toggle between modeling the state of the world and relying on shallow heuristics, depending on which circuit minimizes loss most efficiently. Thus, while the Othello-GPT findings suggest that LLMs can maintain an internal representation of the state of the external world, they also warn us that they will not necessarily do so when a shortcut is available.
@@ 4.1 A Simple Circuit: The Induction Head
A simple example of a circuit is the induction head, first identified in a two-layer toy model by Elhage et al., (2021). This circuit provides a mechanism for a basic but crucial form of in-context learning: completing repeating patterns. Specifically, it enables a model to solve tasks of the form “[A] [B] … [A] → [B].” This is an extremely useful pattern for natural language, as it allows the model to correctly complete a name like “Michael Jordan” by recognizing that this instance of “Michael” is the one that was previously followed by “Jordan” in the same text, rather than “Michael Jackson” or “Michael Faraday.”
@@
  • Step 2: Retrieving the Completion. Later in the model, a second attention head—the induction head proper—processes the second instance of token [A]. Its query essentially asks: “Is there a token earlier in the text that was marked as ‘coming after a token like me’?” It scans the keys of all previous tokens and finds a strong match with the first token [B], which was marked by the prefix head. The induction head then copies the relevant information from [B]’s residual stream into the current stream, strongly upweighting the logit for [B] as the next token. [^26] Remember from section 2 that the information gating is governed by key-query matching. The first attention head at layer one does this matching using positional encodings: it matches a query “I’m looking for position x-1” with the key “I’m at position x-1.” The second attention head at layer two matches the query “I’m looking for a token after an [A]” with the key “I’m after an [A].” This allows information to flow in the desired way. For a great visual explanation, we recommend (McDougall,, 2023).
While this learned circuit still falls short of embodying a principle connecting facts, it constitutes an important first step towards it, because it implements a content-agnostic algorithm. [^27] On the importance of non-content-specific representations in neural networks, see Shea, (2021). The mechanism works just as well for completing “Harry Potter” as it does for “Marie Curie” or for repeating arbitrary sequences of random tokens. It is not learning a content-specific statistical association between “Michael” and “Jordan”; rather, it is implementing a general procedure for sequence completion by looking back in a targeted way.
The initial evidence for induction heads came from toy models. But subsequent behavioral analyses suggest that they are also a crucial component of in-context learning in large-scale models “in the wild” (Olsson et al.,, 2022). [^28] They also study induction heads responsible for “fuzzy” copying tasks of the form [A*] [B*] … [A] → [B], such as in translation where [A*] and [B*] correspond to [A] and [B] in another language. Along those lines, Feucht et al. (2025) recently proposed and empirically investigated a distinction between token induction heads and concept induction heads. While their evidence is primarily behavioral (see Variengien,, 2023), it clearly emerges that models are able to look back in the text in targeted ways to recall relevant token sequences. The induction head thus serves as a compelling first piece of evidence that LLMs can unify many specific instances of pattern completion under a single, learned circuit.
@@ 4.2 A Deeper Principle: The Fourier Algorithm for Modular Addition
While the induction head circuit does not connect facts or embody principled understanding, MI research has also found a striking example of a circuit that does seem to embody principled understanding. Nanda et al. () trained a simple one-layer transformer exclusively on examples of modular addition. Modular addition is addition on a circle or loop instead of a straight number line. This ensures that the numbers are always kept within a specific range, as they are with the hours on a clock or the days of the week. To perform modular addition, one adds the numbers normally before applying the modulus (the size of the loop): one calculates how many times the modulus fits into the sum of the numbers and takes the remainder. This remainder is the solution. When adding 9 + 5 with a modulus of 12, for example, one first calculates 9 + 5 = 14 before seeing how many times 12 goes into 14. Since 12 goes into 14 only once and leaves a remainder of 2, the solution to 9 + 5 (modulo 12) is 2. The model was trained on examples of modular addition using a modulus of 113 (e.g. 55 + 71 = 13), but was never given the rules of addition or the definition of the modulo operator. If LLMs were simply n-gram models, one would expect the model to approach this task by simply memorizing the entire addition table given in the training data.
@@
This simple circuit functionally satisfies the main conditions for explanatory understanding discussed in contemporary epistemology (see Baumberger et al., 2017). It embodies a network of explanatory connections (Khalifa,, 2017; Strevens,, 2013), which satisfy the condition of factivity (Kvanvig,, 2009; Kelp,, 2015), since they rely on geometric facts (the trigonometric identities) that actually obtain. These connections enable the unification of many mathematical facts under a single principle (Friedman,, 1974; Kitcher,, 1981), which in turn allows the model to compress its repository of memorized cases down to a single circuit implementing this principle (Wilkenfeld,, 2018). This also bestows upon the model the inferential capacities associated with grasping (Hills,, 2015; Grimm,, 2011)—manifested here in the model’s ability to generalize correctly to all inputs, not just those in the training set. Finally, the circuit satisfies multiple criteria linked to the condition of justification: it has a wide scope, is used consistently and reflects a pursuit of simplicity via weight decay (Baumberger et al.,, 2017; Beckmann,, 2025).
@@ 4.3 Circuits in the Wild
While the modular addition circuit was found in a toy model, a growing body of research is uncovering similarly complex circuits in state-of-the-art LLMs. Research on regular (non-modular) addition “in the wild” suggests that real LLMs also rely on operations across multiple frequencies—or, equivalently, across multiple moduli (Zhou et al.,, 2024; Ameisen et al.,, 2025; Kantamneni and Tegmark,, 2025; Nikankin et al.,, 2025). Other examples of circuits in the wild include an indirect object identification circuit (Wang et al.,, 2023), a greater-than circuit for numerical comparison (Hanna et al.,, 2023), and a multiple-choice question circuit (Lieberum et al.,, 2023).
Recently, researchers at Anthropic have developed techniques to automatically extract functional circuits for specific prompts from frontier models like Claude 3.5 Sonnet (Ameisen et al.,, 2025; Lindsey et al.,, 2025). They have successfully identified circuits responsible for tasks as diverse as entity recognition, planning the structure of a poem, generating medical diagnoses, and even refusing to answer unsafe queries. A particularly elegant example is a line-breaking circuit in Claude 3.5 Haiku that uses geometric transformations—rotating representations of character count relative to line length in its attention blocks—to detect when the model is approaching the end of a line (Gurnee et al.,, 2025). These findings confirm that the principle of computation via specialized circuits is not an artifact of toy models, but a core operating principle of modern LLMs.
@@
The primary testing ground for this capability is the Abstraction and Reasoning Corpus (ARC-AGI), which was expressly designed to test a system’s ability to discover the rules or principles underlying something at inference time, i.e. after training has been completed (Chollet,, 2019; Chollet et al.,, 2025). To succeed on ARC, a model must be able to look at a few pairs of inputs and outputs, infer the abstract rule by which one is transformed into the other (e.g., “gravity pulls all blue pixels down until they hit a red pixel”), and apply that rule to a test input.
While humans find these tasks intuitive, LLMs typically have a harder time discovering principles in real time. [^31] A related limitation emerges from recent investigations into LLMs’ maze-solving capabilities (Palod et al.,, 2025; Valmeekam et al.,, 2025). This tendency to default to static, training-induced mechanisms has led critics to argue that LLMs lack true intelligence, reducing them to combinators of “heuristics” (Mitchell,, 2025), “program templates” (Chollet,, 2024), or pieces of “procedural knowledge” (Kambhampati,, 2025).
@@ 5  A Motley Mix of Mechanisms
But does the existence of these capacities mean that LLMs actually understand in typical cases? Zooming out from the most compelling, fine-grained findings of mechanistic interpretability to date reveals a more complex reality. Insofar as LLMs instantiate forms of cognition, they instantiate highly strange forms of it, shaped by constraints and incentives that differ starkly from those shaping human intelligence (Fleuret,, 2023; Shanahan, Murray,, 2024). This divergence complicates the attribution of understanding.
@@
At inference, these mechanisms operate in parallel: LLMs typically tackle problems by deploying multiple distinct mechanisms simultaneously rather than relying on a single circuit (Lindsey et al.,, 2025). These mechanisms may cooperate, compete, or contribute independent pieces, with the final output emerging from their combined effect. Higher-quality circuits are sometimes reinforced, but sometimes also drowned out by lower-quality circuits.
@@
The constructive interference observed in modular addition offers a glimpse of this: the model’s final answer arises from the aggregation of five separate frequency-based calculations. Similarly, internal circuitry for standard addition exemplifies the deployment of numerous independent heuristics, such as multiple modulo mechanisms—e.g., computing a + b modulo 2 to determine if the sum is even or odd, and modulo 10 to determine its final digit (Zhou et al.,, 2024; Ameisen et al.,, 2025; Kantamneni and Tegmark,, 2025; Nikankin et al.,, 2025). These are supplemented by parallel range-finding mechanisms of varying precision, which determine the approximate magnitude of the sum (Ameisen et al.,, 2025; Nikankin et al.,, 2025). The final answer emerges from the collective consensus of this “bag of tricks.” [^32] To compute 36 + 59, the LLM Claude 3.5 relies on the additive combination of a range-finding mechanism placing the sum at $\sim 92$; a modulo 100 mechanism according to which the sum ends with 95; and a modulo 10 mechanism according to which the sum ends with 5. Together, these mechanisms constructively interfere at 36 + 59 = 95 (Ameisen et al.,, 2025). Interestingly, the modulo mechanisms resemble how humans carry over digits when adding—effectively adding in steps of modulo 10, 100, 1000, and so on.
@@
Othello-GPT likewise deploys a mix of mechanisms in parallel: localized decision rules—like one that updates specific squares only when A4 is played with certain neighbors occupied; specific board pattern detectors that look for states that make moves legal; and timing-based heuristics that rely on “clock neurons” to track the game’s progress and upweight moves that are statistically more likely in the late game, such as those targeting edge squares (Lin et al.,, 2024).
@@
Even factual recall relies on distinct pathways operating in parallel. When prompted with “Michael Jordan plays the sport of,” Claude 3.5 uses at least two pathways. One pathway, activated by words like “plays” and “sport,” broadly boosts sports-related features, including both “basketball” and “football.” Simultaneously, a second pathway, triggered by “Michael Jordan,” strongly activates “basketball” while suppressing alternatives. The correct output—“basketball”—emerges from the confluence of these separate streams (Ameisen et al.,, 2025; Chughtai et al.,, 2024).
@@
The strangeness of LLM cognition is compounded by the drastic variation in the quality of these mechanisms. They form a “motley mix” spanning the full hierarchy of mechanisms: from pure Blockhead-style memorization of training data (Carlini et al.,, 2023; Aerni et al.,, 2024; Stoehr et al.,, 2024) through shortcuts that rely on spurious correlations (Geirhos et al.,, 2020) and shallow heuristics that rely on superficial statistics (e.g., Ma et al., 2025) to mechanisms encoding the kinds of conceptual, state-of-the-world, and principled understanding we have detailed.
@@
Beyond examining particular examples, this motley mix phenomenon can be established systematically through loss curvature analysis (Merullo et al.,, 2025). This approach asks a simple question: which parts of a model are such that even small weight changes cause large increases in average loss across different inputs, and which parts are such that even large changes barely affect average loss? This can be measured by examining the curvature of the loss landscape with respect to weights, averaged across many inputs. Weights implementing mechanisms recruited across many inputs reinforce the increase in loss and appear as high-curvature; weights recruited only in specific cases produce sharp increases in loss only in those cases, which wash out when averaged, registering as low-curvature or “flat.”
@@
An LLM is thus a motley mix of mechanisms: a system where many circuits of varying quality crystallize during training and operate in parallel at inference. These mechanisms occasionally interact via activation or inhibition, but seem to lack centralized or hierarchical control. [^33] Adding a form of hierarchical control to transformer-based models is currently a hot topic in AI (Jolicoeur-Martineau,, 2025; Behrouz et al.,, 2025). Note that the meta-cognitive features discussed in subsection 2.1 represent early signs of hierarchical processing, but function more as scattered and isolated checks. LLM outputs accordingly resemble the verdict of a gigantic committee of drastically varying expertise, working largely in isolation with only occasional collaboration.